Demo file for DVIWIN 2.7 June 15, 1993 Hippocrates Sendoukas

This driver lets you preview and print DVI files under MS-Windows 3.1. Its main advantages are: speed, compatibility with any raster device with a Windows driver, and graphics capability. Apart from some assembly acceleration for all machines, it uses special routines on 80386 or 80486 processors to speed up the rasterization. All screen and printer handling is done through Windows, so it should work on any printer supported by the system. You can insert arbitrary graphics files produced by most Windows applications, or other standard graphics files (eg. TIFF, PCX, etc.) provided that you have the appropriate graphics filter. The following graph was made with Excel:

Sample Figure: Utility surface for two goods

For more information on using the programs, please read or print the files “dviwin.wri”, “wbr.wri” and “clipmeta.wri”. If you encounter any installation or operation problems, please read the file “helpme.wri” and if it does not resolve the problem, please let me know. My e–mail address is “sendouk@scf.usc.edu”.